OpenStudy (anonymous):

Prove the statement by mathematical induction. 3 + 5 + 7 + . . . + (2n + 1) = n(n + 2)

Parth (parthkohli):

@terenzreignz Would be use \(1 + 3 + 5 + 7\cdots (2n-1) = n^2\)?

Parth (parthkohli):

we*

terenzreignz (terenzreignz):

Why ask me? :D I'm just waiting for you... we might have different approaches :) (NOTE: Whatever approach you're using, it's not the same as what I did ^.^ )

terenzreignz (terenzreignz):

Besides, instructions clearly say "use induction" so I did just that :D

Parth (parthkohli):

OK, write down your proof lol

terenzreignz (terenzreignz):

Oh you -.-

terenzreignz (terenzreignz):

Mathematical Induction... BASE STEP Show it holds for n = 1... Well, clearly... 2n + 1, when n = 1 = 3 =n(n+2) when n = 1 =1(1+2) = 3

OpenStudy (anonymous):

Thank you both for helping by the way!! :)

terenzreignz (terenzreignz):

Okay... INDUCTION STEP You assume it holds for n = k Then you try to show that it holds for n = k+1 Can you do that, @theequestrian ? :) And that goes for you too... @ParthKohli -.-

OpenStudy (anonymous):

Super quick question. How do you know n=1 & n-k & n=k+1?

terenzreignz (terenzreignz):

Those are the usual steps for proving by Mathematical Induction Step 1, show that it holds true for n = 1 Step 2, ASSUME it holds true for n = k, and then show it holds for n = k+1

OpenStudy (anonymous):

Ohh. Okay. So do you just substitute k for n?

terenzreignz (terenzreignz):

Having said that... we assume, therefore, that it holds for n = k In other words, we assume that this equation is true... \[\large 3+5+...(2k+1)=k(k+2)\]

terenzreignz (terenzreignz):

Yes, @theequestrian :)

Parth (parthkohli):

Oh lol, I got confused with that equation

Parth (parthkohli):

Humans make mistake. I've become a little too much of a human.

OpenStudy (anonymous):

Okay, well thats simple enough so far! Haha so our second step would be.. 3 + 5 + 7 + ... + (2k + 1) = k(k + 2)

Parth (parthkohli):

mistakes* ^ another one -_-

Parth (parthkohli):

Wouldn't it be proving that\[3 + 5 + 7 + \cdots 2k+1 + 2(k+1) + 1 = (k+1)(k+3)\]

terenzreignz (terenzreignz):

Is that so bad, @ParthKohli ? :P We assume that this \[\large 3+5+...(2k+1)=k(k+2)\]is true, and test this out... \[\large 3+5+...(2k+1)\color{green}{+[2(k+1)+1]}=(k\color{green}{+1)}[(k\color{green}{+1})+2]\]

terenzreignz (terenzreignz):

Understood, @theequestrian ? @ParthKohli You may have a different way of using PMI... do you assume it holds for n = k-1 and show that it holds for n = k instead? It looks to me like you're getting things mixed up...

Parth (parthkohli):

Yeah, I'm getting them mixed up -_-

OpenStudy (anonymous):

I'm actually a little lost..

Parth (parthkohli):

I quit. :-P

terenzreignz (terenzreignz):

Okay, let's get you up to speed... we assume it is true for n = k, am I right? So we replace n with k... \[\large 3+5+...(2k+1)=k(k+2)\]

OpenStudy (anonymous):

Okay, I get that. Thats about all I get.. Lol & thank you!

terenzreignz (terenzreignz):

Okay. Next we have to show that it holds for n = k+1... so you see this? \[\large 3+5+...(2\color{red}n+1)=\color{red}n(\color{red}n+2)\] I just replaced all instances of n with a k+1, like so... \[\large 3+5+...[2\color{red}{(k+1)}+1]=\color{red}{(k+1)}[\color{red}{(k+1)}+2]\] Now do you understand? :)

OpenStudy (anonymous):

Ohhh. So you took the original equation and substituted for k & then went BACK to the original and substituted for (k+1)? What was the point of putting k in alone?

terenzreignz (terenzreignz):

The point, we assumed that the one involving k (alone) is already true, and now, all we have to show (prove) is that the one involving k+1 is ALSO true, given that the one involving k (alone) is true... get it?

OpenStudy (anonymous):

Yes, thank you! & I have to fill in some blanks & I've done some of them, can you check them?

terenzreignz (terenzreignz):

What? Is this the same question?

OpenStudy (anonymous):

Yes it is

terenzreignz (terenzreignz):

All right...

OpenStudy (anonymous):

I was trying to figure it out & then fill in the blanks, but its actually harder than I had planned

terenzreignz (terenzreignz):

for slot a3, you should know better... remember, it's just the original equation you were trying to prove, save for the fact that we replaced all instances of n with k...

OpenStudy (anonymous):

Ohh, so for that line its just simply replacing all the n's with k's?

OpenStudy (anonymous):

Oh duh because it says k=n

terenzreignz (terenzreignz):

Yes. -.-

OpenStudy (anonymous):

So would a4 = 2 a5 = k a6= 2 a7 = 1 I think?

terenzreignz (terenzreignz):

a5? no... look at this more carefully... this is the n=k+1 part, right? \[\large 3+5+...\boxed{[2\color{red}{(k+1)}+1]}=\color{red}{(k+1)}[\color{red}{(k+1)}+2]\]

terenzreignz (terenzreignz):

It's [2(k+1) + 1] right? Mind simplifying that for me? :P

OpenStudy (anonymous):

2k+ 2 +1 2k+3

OpenStudy (anonymous):

2k+3 is my answer haha

terenzreignz (terenzreignz):

Then, a5 should instead be...?

OpenStudy (anonymous):

it should be 3

terenzreignz (terenzreignz):

Yes :) Note that a6 and a7 basically mirror a4 and a5

OpenStudy (anonymous):

Oh okay, so it would be a6 = 2 a7 = 3

terenzreignz (terenzreignz):

Yup :)

OpenStudy (anonymous):

Thank you soo much!! I kind of get it now:) Couldn't say that before! I really appreciate all your help! Thanks!

terenzreignz (terenzreignz):

No problem :)